# Low-energy critical behavior in two-dimensional tilted semi-Dirac semimetals driven by fermion-fermion interactions

### v3

@@ Abstract
Employing the renormalization group approach, we carefully investigate the critical behavior of two-dimensional tilted semi-Dirac semimetals induced by the fermion-fermion interactions in the low-energy regime. After incorporating all one-loop corrections, we derive the coupled RG equations of all related parameters and introduce two distinct strategies, named as Strategy I and Strategy II, to describe different scenarios. A detailed numerical analysis yields several interesting behavior in the low-energy limit. At first, we notice that the fermion-fermion interactions either vanish or diverge in the Strategy I, depending on the initial values of the tilting parameter and the fermionic couplings, whereas these interactions in the Strategy II always diverge at a certain critical energy scale, which is associated with the initial conditions. Next, the microstructural parameter $\alpha$ and the fermion velocity $v_{F}$ in the Strategy I share the similar behavior with their Strategy II counterparts. It is observed that fermion-fermion interactions lead to an increase in $\alpha$ while driving a decrease in $v_{F}$. Furthermore, the system can either be attracted by the Gaussian fixed point (GFP) or certain relatively fixed point (RFP) in the Strategy I. However, it always flow towards the RFP in the Strategy II at the lowest-energy limit. These results would provide helpful insights into the studies on observable quantities and phase transitions in the two-dimensional tilted semi-Dirac semimetals and the analogous semimetals.
@@ I Introduction
In principle, the 2D tilted SDSMs can be categorized into two scenarios heavily depending on the strength of tilting parameters Neto2009RMP ; Peres2010RMP ; Jafari ; Lee2018PRB ; Jafari2019PRB-t ; Lee2019PRB ; Soluyanov2015Nature ; Noh2017PRL ; Fei2017PRB ; Yan2017NC , namely, type-I tilted SDSMs featured by intact Dirac points (cones) Neto2009RMP ; Peres2010RMP ; Jafari and type-II tilted SDSM characterized by the open Fermi surface (such as two straight lines, etc.)  Soluyanov2015Nature ; Noh2017PRL ; Fei2017PRB ; Yan2017NC . Without loss of generality, we put our focus on the type-I situation within this work. In this circumstance, it is of significance to investigate the critical behavior of such type-I tilted SDSMs, which is caused by the unconventional low-energy excitations accompanied by unique Dirac cones in tandem with their intimate interactions in the low-energy regime. Regarding the interactions in Dirac materials, two types are particularly important: Coulomb interactions and fermion-fermion interactions, which may play an important role in pinning down their low-energy properties. In principle, the long-range Coulomb interaction can be easily screened in the realistic systems via adding certain substrate to realize a large dielectric constant Neto2009RMP ; Kotov2012RMP ; Sarma2011RMP ; Katsnelson2006PRB . As a result, its effects on the low-energy properties of tilted SDSMs are more important  Lee2018PRB ; Lee2019PRB ; Fritz2017PRB ; Fritz2019arXiv . In comparison, the fermion-fermion interactions can still be intact and henceforth they are closely associated with the unusual signatures in fermionic systems BXW2023EPJP ; DZZW2020PRB ; Kivelson2009PRL ; Murray2014PRB ; Herbut ; Herbut-2 ; Wang2017PRB_QBCP ; Wang2018 ; Fradkin2010ARCMP ; Roy2004arXiv ; Nandkishore2012NP ; Wang-NPhyB ; Sur2016NJP ; Roy2016JHEP ; Han2017PRB ; Roy2018PRX2 ; Mandal2018PRB ; Sur2019PRL ; Szabo2021 ; Wang2023EPJB ; Fu2023arXiv while the Coulomb interaction is screened Katsnelson2006PRB ; Neto2009RMP ; Kotov2012RMP ; Sarma2011RMP . However, as to the 2D tilted SDSMs, it is worth highlighting that the fermion-fermion interactions have hitherto been paid relatively less attention, which may lead to the neglect or omission of valuable physical information. Accordingly, it is therefore of particular significance to systematically examine the role of fermion-fermion interactions in determining the low-energy fates of 2D tilted SDSMs. Unambiguously elucidating this issue would be great helpful to reveal the unique properties of 2D semi-Dirac materials and even instructive to explore new Dirac-like materials.
@@
At first, we find that the fermion-fermion interactions tend to either vanish or diverge in the low-energy regime. In the Strategy I, the bigger initial values of the tilting parameter and the fermionic couplings cause fermion-fermion interactions to diverge at the low-energy limit. On the contrary, for smaller initial values, the fermion-fermion interactions diminish and eventually vanish as energy decreases. In comparison, as for the Strategy II, the fermion-fermion interactions always diverge at certain critical energy scale labeled by $l_{c}$, regardless of the initial parameter values. However, for both strategies, the critical energy scale can be increased by the bigger initial values of fermionic interactions ($\lambda_{i0}$) and the tilting parameter $(\zeta)$, in tandem with the smaller values of the fermion velocity $(v_{F0})$. Besides, a smaller microstructural parameter $(\alpha_{0})$ increases the the critical energy scale in the Strategy I but instead decreases in the Strategy II.
@@ II Microscopic model and effective theory
The noninteracting Hamiltonian of a 2D SDSM with Dirac points takes the form of Saha2016PRB ; Montambaux
@@
where $\alpha$ is a microstructural parameter that measures the inverse of quasiparticle mass along $x$, and $v_{F}$ represents the Fermi velocity along the $y$ direction, respectively. Hereby, $\sigma_{0}$ denotes the $2\times 2$ identity matrix, and $\sigma_{i}$ with $i=1,2,3$ characterize the Pauli matrices, which satisfy the algebra $\{\sigma_{i},\sigma_{j}\}=2\delta_{ij}$.
@@ IV Two distinct strategies for the parameter $\eta$
As aforementioned in Sec. III, a scaling parameter $\eta$ has been introduced to construct the RG transformations that carry the features of a 2D tilted SDSM. Although the exact value of $\eta$ is unclear, we stop hereby and turn our attention to the potential expressions of this very parameter. Due to the unique dispersion of the 2D tilted SDSMs, $\eta$ is confined within the range $\eta\in(-1/2,0)$, with the boundaries $\eta=0$ and $\eta=-1/2$ corresponding to the linear Dirac SMs Neto2009RMP and the QBCP SMs Kivelson2009PRL ; Wang2017PRB_QBCP ; Murray2014PRB ; DZZW2020PRB ; Fu2023arXiv , respectively. To simplify our analysis, it is helpful to consider two distinct situations including Strategy I with a constant $\eta$ and Strategy II where $\eta$ is also dependent upon other parameters, i.e., a parameter-dependent $\eta$.
@@ V.1 Behavior of the strengths of fermion-fermions interactions
We begin with the Strategy I. Our numerical analysis suggests that whether fermion-fermion interactions diverge in the low-energy regime is closely dependent on the initial conditions. When the initial values of the fermion-fermion interactions and the tilting parameter are small, Fig. 2(a) with $\zeta=0.3$ shows that fermionic couplings decrease as the energy scale decreases (or $l$ increases), eventually vanishing in the low-energy regime. Besides, this qualitative results are insensitive to the variations of both the fermion velocity ($v_{F0}$) and the microstructural parameter ($\alpha_{0}$). In sharp contrast, while the initial values of fermionic couplings are taken big enough, the fermion-fermion interactions ($\lambda_{i}$ with $i=0,1,2,3$) as illustrated in Fig. 2(b) with $\zeta=0.5$ quickly increase and then finally diverge at certain critical energy scale denoted by $l_{c}$. Hereby, it is of particular importance to emphasize that the critical energy scale $l_{c}$ is closely associated with the initial values of parameters $\zeta$, $v_{F}$, $\alpha$, and $\lambda_{i}$. Let us take $\lambda_{0}$ for example (the basic tendencies for $\lambda_{1,2,3}$ are analogous). In particular, a bigger tilting parameter $\zeta$ is helpful to the divergence of fermionic couplings as shown in Fig. 3(a). Additionally, Fig. 4(a)-(c) indicate a smaller $v_{F0}$ and $\alpha_{0}$ as well as a bigger $\lambda_{i0}$ can enhance the divergence and diminish the $l_{c}$.
@@ V.2 Fate of the microstructural parameter $\alpha$ and the Fermi velocity $v_{F}$
Specifically, let us begin with the Strategy I. On one side, if the fermion-fermion couplings are not divergent as shown in Fig. 5(a) and Fig. 6(a) with the small initial values and tilting parameter, $\alpha/\alpha_{0}$ monotonously increases, while $v_{F}/v_{F0}$ gradually decreases and vanishes with reducing the energy scale. On the other side, once the fermion-fermion interactions, armed with the bigger $\lambda_{i0}$ and $\zeta$, go towards divergence at a critical energy scale denoted by $l_{c}$, the RG equations are required to stop at $l=l_{c}$. This implies that the $\alpha$ increases and reaches its maximum achievable value $\alpha(l_{c})$ as depicted in Fig. 5(b). Meanwhile, as illustrated in Fig. 6(b), the Fermi velocity falls until it arrives at a finite value, i.e., $v_{F}(l_{c})$. Besides, the basic evolutions of both $\alpha/\alpha_{0}$ and $v_{F}/v_{F0}$ are insensitive to the tilting parameter, $v_{F0}$, $\alpha_{0}$, and $\lambda_{i0}$.
@@
To proceed, we consider the Strategy II. Learning from Fig. 5(c)-(f) and Fig. 6(c)-(f), we notice that the basic properties of both $\alpha/\alpha_{0}$ and $v_{F}/v_{F0}$ analogous to their Strategy I counterparts. To be specific, Fig. 5(c)-(d) and Fig. 6(c)-(d) suggest that the tendencies of $\alpha$ or $v_{F}$ in Strategy II are insusceptible to the tilting parameter. In contrast to the Strategy I, $v_{F0}$ and $\alpha_{0}$ are capable of considerably reshape the slopes of $\alpha/\alpha_{0}$ or $v_{F}/v_{F0}$ as illustrated in Fig. 5(e)-(f) and Fig. 6(e)-(f). Moreover, we realize that the critical energy scales are closely associated with the initial values of interaction parameters. In particular, Fig. 5(e)-(f) and Fig. 6(e)-(f) indicate that a smaller $\alpha_{0}$ yields a lower critical energy scale (a bigger $l_{c}$), but instead a smaller initial value of Fermi velocity gives rise to a higher critical energy scale (a smaller $l_{c}$).
@@ VI Summary
At first, we notice that the behavior of fermion-fermion interactions in the low-energy regime can either vanish or diverge. With respect to the Strategy I, the low-energy behavior of fermion-fermion interactions depends on both the interaction strength itself and the initial value of the tilting parameter as shown in Fig. 2. For the weak initial values of the interactions and tilting parameter, the fermion-fermion couplings decrease and eventually vanish. In contrast, if the initial values of the related parameter are sufficiently large, they can increase and diverge at $l=l_{c}$. However, as for the Strategy II, regardless of the initial values of the relevant parameters, the fermion-fermion couplings can diverge at a certain critical energy scale. In particular, the bigger tilting parameter, the microstructural parameter, and fermion-fermion interactions can increase the critical energy scale, corresponding to a smaller $l_{c}$, but instead the bigger fermion velocity prefers to decrease the critical energy scale. Next, we examine the low-energy behavior of the microstructural parameter $\alpha$ and the fermionic velocity $v_{F}$. It can be found that the fermion-fermion couplings drive an increase in $\alpha$ but a decrease in $v_{F}$, respectively. As to the Strategy I, if the fermion-fermion interactions does not diverge, $\alpha$ increases and $v_{F}$ decreases gradually in the low-energy regime. In contrast, once the fermion-fermion interactions $\lambda_{i}$ diverge at the critical energy $l_{c}$ for either the Strategy I or the Strategy II, the microstructural parameter and fermion velocity would be truncated at $l=l_{c}$, namely obtaining the final values $\alpha(l_{c})$ and $v_{F}(l_{c})$ as listed in Table. 1 for distinct kinds of initial conditions. Furthermore, we find that, depending the fermion-fermion interactions vanish or diverge in the low-energy regime, the system can either flow towards the GFP or RFP, which is defined by $\lambda_{j}(l_{c})/\lambda_{0}(l_{c})$, as schematically shown in Fig. 9. With a fixed tilting parameter, the RFP depicted in Fig. 7 is very stable against the variations of other parameters. However, Fig. 8 indicates that the RFP is sensitive to the initial value of tilting parameter $\zeta$ for both strategies. To recapitulate, we show that the fermion-fermion interactions are closely associated with the related parameters in the low-energy regime, which induce certain RFP at the critical energy scales. Future research are necessary to explore the observable quantities and phase transitions for the 2D tilted SDSM and analogous materials. These may offer valuable insights and scientific significance for future applications and related studies.

### v4

@@ Abstract
Employing the renormalization group approach, we carefully investigate the critical behavior of two-dimensional tilted semi-Dirac semimetals induced by the fermion-fermion interactions in the low-energy regime. After incorporating all one-loop corrections, we derive the coupled RG equations of all related parameters and introduce two distinct strategies, named as Strategy I and Strategy II, to describe different scenarios. A detailed numerical analysis yields several interesting behavior in the low-energy limit. At first, we notice that the fermion-fermion interactions either vanish or diverge in the Strategy I, depending on the initial values of the tilting parameter and the fermionic couplings, whereas these interactions in the Strategy II always diverge at a certain critical energy scale, which is associated with the initial conditions. Next, the microstructural parameter $\alpha$ and the fermion velocity $v_{F}$ in the Strategy I share the similar behavior with their Strategy II counterparts. It is observed that fermion-fermion interactions lead to an increase in $\alpha$ while driving a decrease in $v_{F}$. Furthermore, the system can either be attracted by the Gaussian fixed point (GFP) or certain relatively fixed point (RFP) in the Strategy I. However, it always flows toward the RFP in the Strategy II at the lowest-energy limit. These results would provide helpful insights into the studies on observable quantities and phase transitions in the two-dimensional tilted semi-Dirac semimetals and the analogous semimetals.
@@ I Introduction
In principle, the 2D tilted SDSMs can be categorized into two scenarios heavily depending on the strength of tilting parameters Neto2009RMP ; Peres2010RMP ; Jafari ; Lee2018PRB ; Jafari2019PRB-t ; Lee2019PRB ; Soluyanov2015Nature ; Noh2017PRL ; Fei2017PRB ; Yan2017NC , namely, type-I tilted SDSMs featured by intact Dirac points (cones) Neto2009RMP ; Peres2010RMP ; Jafari and type-II tilted SDSM characterized by the open Fermi surface (such as two straight lines, etc.)  Soluyanov2015Nature ; Noh2017PRL ; Fei2017PRB ; Yan2017NC . Without loss of generality, we put our focus on the type-I situation within this work. In this circumstance, it is of significance to investigate the critical behavior of such type-I tilted SDSMs, which is caused by the unconventional low-energy excitations accompanied by unique Dirac cones in tandem with their intimate interactions in the low-energy regime. Regarding the interactions in Dirac materials, two types are particularly important: Coulomb interactions and fermion-fermion interactions, which may play an important role in pinning down their low-energy properties. In principle, the long-range Coulomb interaction can be easily screened in the realistic systems via adding certain substrate to realize a large dielectric constant Neto2009RMP ; Kotov2012RMP ; Sarma2011RMP ; Katsnelson2006PRB . As a result, its effects on the low-energy properties of tilted SDSMs are more important  Lee2018PRB ; Lee2019PRB ; Fritz2017PRB ; Fritz2019arXiv . In comparison, the fermion-fermion interactions can still be intact, and henceforth, they are closely associated with the unusual signatures in fermionic systems BXW2023EPJP ; DZZW2020PRB ; Kivelson2009PRL ; Murray2014PRB ; Herbut ; Herbut-2 ; Wang2017PRB_QBCP ; Wang2018 ; Fradkin2010ARCMP ; Roy2004arXiv ; Nandkishore2012NP ; Wang-NPhyB ; Sur2016NJP ; Roy2016JHEP ; Han2017PRB ; Roy2018PRX2 ; Mandal2018PRB ; Sur2019PRL ; Szabo2021 ; Wang2023EPJB ; Fu2023arXiv while the Coulomb interaction is screened Katsnelson2006PRB ; Neto2009RMP ; Kotov2012RMP ; Sarma2011RMP . However, as to the 2D tilted SDSMs, it is worth highlighting that the fermion-fermion interactions have hitherto been paid relatively less attention, which may lead to the neglect or omission of valuable physical information. Accordingly, it is therefore of particular significance to systematically examine the role of fermion-fermion interactions in determining the low-energy fates of 2D tilted SDSMs. Unambiguously elucidating this issue would be greatly helpful to reveal the unique properties of 2D semi-Dirac materials and even instructive to explore new Dirac-like materials.
@@
At first, we find that the fermion-fermion interactions tend to either vanish or diverge in the low-energy regime. In the Strategy I, the bigger initial values of the tilting parameter and the fermionic couplings cause fermion-fermion interactions to diverge at the low-energy limit. On the contrary, for smaller initial values, the fermion-fermion interactions diminish and eventually vanish as energy decreases. In comparison, as for the Strategy II, the fermion-fermion interactions always diverge at certain critical energy scale labeled by $l_{c}$, regardless of the initial parameter values. However, for both strategies, the critical energy scale can be increased by the bigger initial values of fermionic interactions ($\lambda_{i0}$) and the tilting parameter $(\zeta)$, in tandem with the smaller values of the fermion velocity $(v_{F0})$. Besides, a smaller microstructural parameter $(\alpha_{0})$ increases the critical energy scale in the Strategy I but instead decreases in the Strategy II.
@@ II Microscopic model and effective theory
The non-interacting Hamiltonian of a 2D SDSM with Dirac points takes the form of Saha2016PRB ; Montambaux
@@
where $\alpha$ is a microstructural parameter that measures the inverse of quasiparticle mass along $x$ and $v_{F}$ represents the Fermi velocity along the $y$ direction, respectively. Hereby, $\sigma_{0}$ denotes the $2\times 2$ identity matrix, and $\sigma_{i}$ with $i=1,2,3$ characterize the Pauli matrices, which satisfy the algebra $\{\sigma_{i},\sigma_{j}\}=2\delta_{ij}$.
@@ IV Two distinct strategies for the parameter $\eta$
As aforementioned in Sec. III, a scaling parameter $\eta$ has been introduced to construct the RG transformations that carry the features of a 2D tilted SDSM. Although the exact value of $\eta$ is unclear, we stop here and turn our attention to the potential expressions of this very parameter. Due to the unique dispersion of the 2D tilted SDSMs, $\eta$ is confined within the range $\eta\in(-1/2,0)$, with the boundaries $\eta=0$ and $\eta=-1/2$ corresponding to the linear Dirac SMs Neto2009RMP and the QBCP SMs Kivelson2009PRL ; Wang2017PRB_QBCP ; Murray2014PRB ; DZZW2020PRB ; Fu2023arXiv , respectively. To simplify our analysis, it is helpful to consider two distinct situations including Strategy I with a constant $\eta$ and Strategy II where $\eta$ is also dependent upon other parameters, i.e., a parameter-dependent $\eta$.
@@ V.1 Behavior of the strengths of fermion-fermions interactions
We begin with the Strategy I. Our numerical analysis suggests that whether fermion-fermion interactions diverge in the low-energy regime is closely dependent on the initial conditions. When the initial values of the fermion-fermion interactions and the tilting parameter are small, Fig. 2(a) with $\zeta=0.3$ shows that fermionic couplings decrease as the energy scale decreases (or $l$ increases), eventually vanishing in the low-energy regime. Besides, these qualitative results are insensitive to the variations of both the fermion velocity ($v_{F0}$) and the microstructural parameter ($\alpha_{0}$). In sharp contrast, while the initial values of fermionic couplings are taken big enough, the fermion-fermion interactions ($\lambda_{i}$ with $i=0,1,2,3$) as illustrated in Fig. 2(b) with $\zeta=0.5$ quickly increase and then finally diverge at certain critical energy scale denoted by $l_{c}$. Hereby, it is of particular importance to emphasize that the critical energy scale $l_{c}$ is closely associated with the initial values of parameters $\zeta$, $v_{F}$, $\alpha$, and $\lambda_{i}$. Let us take $\lambda_{0}$ for example (the basic tendencies for $\lambda_{1,2,3}$ are analogous). In particular, a bigger tilting parameter $\zeta$ is helpful to the divergence of fermionic couplings as shown in Fig. 3(a). Additionally, Fig. 4(a)-(c) indicate a smaller $v_{F0}$ and $\alpha_{0}$ as well as a bigger $\lambda_{i0}$ can enhance the divergence and diminish the $l_{c}$.
@@ V.2 Fate of the microstructural parameter $\alpha$ and the Fermi velocity $v_{F}$
Specifically, let us begin with the Strategy I. On one side, if the fermion-fermion couplings are not divergent as shown in Fig. 5(a) and Fig. 6(a) with the small initial values and tilting parameter, $\alpha/\alpha_{0}$ monotonously increases, while $v_{F}/v_{F0}$ gradually decreases and vanishes with reducing the energy scale. On the other side, once the fermion-fermion interactions, armed with the bigger $\lambda_{i0}$ and $\zeta$, go toward divergence at a critical energy scale denoted by $l_{c}$, the RG equations are required to stop at $l=l_{c}$. This implies that the $\alpha$ increases and reaches its maximum achievable value $\alpha(l_{c})$ as depicted in Fig. 5(b). Meanwhile, as illustrated in Fig. 6(b), the Fermi velocity falls until it arrives at a finite value, i.e., $v_{F}(l_{c})$. Besides, the basic evolutions of both $\alpha/\alpha_{0}$ and $v_{F}/v_{F0}$ are insensitive to the tilting parameter, $v_{F0}$, $\alpha_{0}$, and $\lambda_{i0}$.
@@
To proceed, we consider the Strategy II. Learning from Fig. 5(c)-(f) and Fig. 6(c)-(f), we notice that the basic properties of both $\alpha/\alpha_{0}$ and $v_{F}/v_{F0}$ analogous to their Strategy I counterparts. To be specific, Fig. 5(c)-(d) and Fig. 6(c)-(d) suggest that the tendencies of $\alpha$ or $v_{F}$ in Strategy II are insusceptible to the tilting parameter. In contrast to the Strategy I, $v_{F0}$ and $\alpha_{0}$ are capable of considerably reshaping the slopes of $\alpha/\alpha_{0}$ or $v_{F}/v_{F0}$ as illustrated in Fig. 5(e)-(f) and Fig. 6(e)-(f). Moreover, we realize that the critical energy scales are closely associated with the initial values of interaction parameters. In particular, Fig. 5(e)-(f) and Fig. 6(e)-(f) indicate that a smaller $\alpha_{0}$ yields a lower critical energy scale (a bigger $l_{c}$), but instead a smaller initial value of Fermi velocity gives rise to a higher critical energy scale (a smaller $l_{c}$).
@@ VI Summary
At first, we notice that the behavior of fermion-fermion interactions in the low-energy regime can either vanish or diverge. With respect to the Strategy I, the low-energy behavior of fermion-fermion interactions depends on both the interaction strength itself and the initial value of the tilting parameter as shown in Fig. 2. For the weak initial values of the interactions and tilting parameter, the fermion-fermion couplings decrease and eventually vanish. In contrast, if the initial values of the related parameter are sufficiently large, they can increase and diverge at $l=l_{c}$. However, as for the Strategy II, regardless of the initial values of the relevant parameters, the fermion-fermion couplings can diverge at a certain critical energy scale. In particular, the bigger tilting parameter, the microstructural parameter, and fermion-fermion interactions can increase the critical energy scale, corresponding to a smaller $l_{c}$, but instead the bigger fermion velocity prefers to decrease the critical energy scale. Next, we examine the low-energy behavior of the microstructural parameter $\alpha$ and the fermionic velocity $v_{F}$. It can be found that the fermion-fermion couplings drive an increase in $\alpha$ but a decrease in $v_{F}$, respectively. As to the Strategy I, if the fermion-fermion interactions do not diverge, $\alpha$ increases and $v_{F}$ decreases gradually in the low-energy regime. In contrast, once the fermion-fermion interactions $\lambda_{i}$ diverge at the critical energy $l_{c}$ for either the Strategy I or the Strategy II, the microstructural parameter and fermion velocity would be truncated at $l=l_{c}$, namely obtaining the final values $\alpha(l_{c})$ and $v_{F}(l_{c})$ as listed in Table. 1 for distinct kinds of initial conditions. Furthermore, we find that, depending the fermion-fermion interactions vanish or diverge in the low-energy regime, the system can either flow toward the GFP or RFP, which is defined by $\lambda_{j}(l_{c})/\lambda_{0}(l_{c})$, as schematically shown in Fig. 9. With a fixed tilting parameter, the RFP depicted in Fig. 7 is very stable against the variations of other parameters. However, Fig. 8 indicates that the RFP is sensitive to the initial value of tilting parameter $\zeta$ for both strategies. To recapitulate, we show that the fermion-fermion interactions are closely associated with the related parameters in the low-energy regime, which induce certain RFP at the critical energy scales. Future research is necessary to explore the observable quantities and phase transitions for the 2D tilted SDSM and analogous materials. These may offer valuable insights and scientific significance for future applications and related studies.